A petroleum company has two different sources of crude oil. The first source provides crude oil that is 45% hydrocarbons, and the second one provides crude oil that is
95% hydrocarbons. In order to obtain 140 gallons of crude oil that is 55% hydrocarbons, how many gallons of crude oil must be used from each of the two sources?

Answer by TimothyLamb(4379)

x + y = 140
45x + 95y = 55*140
---
put the system of linear equations into standard form
---
x + y = 140
45x + 95y = 7700
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = first source = 112 gal
y = second source = 28 gal
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php